# A Multi-Fidelity Control Variate Approach for Policy Gradient Estimation

### v3

@@ 7 Conclusions
We present a multi-fidelity policy gradient (MFPG) framework, which mixes a small amount of potentially expensive high-fidelity data with a larger volume of cheap lower-fidelity data to construct unbiased, variance-reduced estimators for on-policy policy gradients. We instantiate this framework through a practical, multi-fidelity variant of the standard REINFORCE algorithm. Theoretically, we show that under standard assumptions, the MFPG estimator guarantees asymptotic convergence of multi-fidelity REINFORCE to locally optimal policies (up to first order optimality conditions) in the target high-fidelity environment of interest. Moreover, whenever the statistical correlation between high- and low-fidelity samples is nontrivial, MFPG REINFORCE achieves faster finite-sample convergence than its single-fidelity counterpart. Empirical evaluation of MFPG in high-fidelity data-scarce regimes across a suite of simulated robotics benchmark tasks with varied multi-fidelity transition dynamics highlights both its efficacy and robustness. With mild to moderate dynamics gaps, MFPG reliably improves median performance over a standard high-fidelity-only baseline and matches the performance of leading multi-fidelity methods, despite its simplicity and minimum tuning overhead. Under large gaps, MFPG demonstrates the greatest robustness among the evaluated multi-fidelity approaches. Furthermore, an additional experiment demonstrates that MFPG can remain effective even in the presence of reward misspecification in the low-fidelity environments.
In summary, the proposed framework offers a novel paradigm for leveraging multi-fidelity data in reinforcement learning, providing promising directions for efficient sim-to-real transfer and principled approaches to managing the trade-off between policy performance and data collection costs. Future research spans several avenues. First, extending the MFPG framework to modern RL algorithms—including actor–critic and off-policy/offline methods—raises new questions about bias–variance trade-offs that merit deeper investigation. Second, developing systematic techniques to enhance correlation between high- and low-fidelity samples without introducing bias is critical for further performance gains. Finally, applying MFPG to real-world robotic systems represents an important step toward realizing its potential in practical sim-to-real transfer.
@@ Acknowledgments
We thank Brett Barkley, Jacob Levy, Haoran Xu, and Mustafa Karabag for their helpful discussions.
Appendix table of contents
@@ Appendix A Complete MFPG REINFORCE Algorithm
Input: Initial policy $\pi_{\theta}$, multi-fidelity environments $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$, batch size for correlated trajectory samples $N^{h}$, batch size for control variate sample mean estimate $N^{l}$.
Output: Trained policy $\pi_{\theta}$.
while maximum high-fidelity environment step not reached do
  Collect $N^{h}$ correlated trajectory samples $\mathcal{D}_{1}=\{\tau_{i}^{h},\tau_{i}^{l}\}_{i=1}^{N^{h}}$. $\triangleright$ See Algorithm 1
  Collect $N^{l}$ additional uncorrelated low-fidelity trajectory samples $\mathcal{D}_{2}=\{\tau_{j}^{l}\}_{j=1}^{N^{l}}$.
  Compute correlated policy gradient random variables using $\mathcal{D}_{1}$:
$X^{\pi_{\theta}}_{\tau^{h},i}=\frac{1}{T}\sum_{t=0}^{T-1}(G^{h}_{i,t}-V_{\phi}(s^{h}_{i,t}))\,\log\pi_{\theta}(a^{h}_{i,t}\mid s^{h}_{i,t}),$
$X^{\pi_{\theta}}_{\tau^{l},i}=\frac{1}{T}\sum_{t=0}^{T-1}(G^{l}_{i,t}-V_{\phi}(s^{l}_{i,t}))\,\log\pi_{\theta}(a^{l}_{i,t}\mid s^{l}_{i,t}).$
  Estimate the optimal control variate coefficient $\hat{c}^{*}$ using $\mathcal{D}_{1}$ and Eq.˜2.
  Estimate the control variate sample mean using $\mathcal{D}_{2}$: $\hat{\mu}^{l}=\frac{1}{N^{l}}\sum_{j=1}^{N^{l}}X^{\pi_{\theta}}_{\tau^{l},j}$.
  Construct multi-fidelity policy gradient estimator: $\mathbb{E}[\nabla_{\theta}Z^{\pi_{\theta}}]\approx\frac{1}{N^{h}}\sum_{i=1}^{N^{h}}\nabla_{\theta}\bigl[X^{\pi_{\theta}}_{\tau^{h},i}+\hat{c}^{*}(X^{\pi_{\theta}}_{\tau^{l},i}-\hat{\mu}^{l})\bigr]$.
  Perform policy optimization.
  Fit the value function network $V_{\phi}(s^{l}_{i,t})$ using high-fidelity samples in $\mathcal{D}_{1}$.
return policy $\pi_{\theta}$
@@
Appendix B Proof of Theorem˜1
@@
Thus, for $\theta_{1},\theta_{2}\in^{d}$, we have for the high-fidelity only policy gradient,
$\left\|\nabla J(\theta_{1})-\nabla J(\theta_{2})\right\|=\Bigg\|\sum_{t=0}^{T}\sum_{\kappa=0}^{T-t}\gamma^{t+\kappa}\int\Big(\underbrace{r^{h}(s^{h}_{t+\kappa},a^{h}_{t+\kappa})\left[\nabla\log\pi_{\theta_{1}}(a_{t}^{h}|s_{t}^{h})-\nabla\log\pi_{\theta_{2}}(a_{t}^{h}|s_{t}^{h})\right]p^{h}_{\theta_{1},0:t+\kappa}}_{A}$
$+\underbrace{r^{h}(s^{h}_{t+\kappa},a^{h}_{t+\kappa})\nabla\log\pi_{\theta_{2}}(a_{t}^{h}|s_{t}^{h})\left(p^{h}_{\theta_{1},0:t+\kappa}-p^{h}_{\theta_{2},0:t+\kappa}\right)}_{B}\Big)ds^{h}_{0:t+\kappa}da^{h}_{0:t+\kappa}\Bigg\|$
$\leq\sum_{t=0}^{\infty}\sum_{\kappa=0}^{\infty}\gamma^{t+\kappa}\Bigg[\underbrace{\int\Big(\left|r^{h}(s^{h}_{t+\kappa},a^{h}_{t+\kappa})\right|\left\|\nabla\log\pi_{\theta_{1}}(a_{t}^{h}|s_{t}^{h})-\nabla\log\pi_{\theta_{2}}(a_{t}^{h}|s_{t}^{h})\right\|p^{h}_{\theta_{1},0:t+\kappa}ds^{h}_{0:t+\kappa}da^{h}_{0:t+\kappa}}_{I_{1}}$
$+\underbrace{\int\left|r^{h}(s^{h}_{t+\kappa},a^{h}_{t+\kappa})\right|\left\|\nabla\log\pi_{\theta_{2}}(a_{t}^{h}|s_{t}^{h})\right\|\left|p^{h}_{\theta_{1},0:t+\kappa}-p^{h}_{\theta_{2},0:t+\kappa}\right|ds^{h}_{0:t+\kappa}da^{h}_{0:t+\kappa}}_{I_{2}}\Bigg]$ (3)
@@
$\sum_{t=0}^{T}\sum_{\kappa=0}^{T-t}(t+\kappa+1)\gamma^{t+\kappa}$ $=\frac{\left(1-(T+2)^{2}\gamma^{T+1}+(T+3)(T+1)\gamma^{T+2}\right)(1-\gamma)+2\gamma\left(1-(T+2)\gamma^{T+1}+(T+1)\gamma^{T+2}\right)}{(1-\gamma)^{3}},$
@@
Under Assumptions˜1, 2, 3 and 4, the norm of the Multi-Fidelity Policy Gradient estimator is bounded, i.e., $\exists~\hat{\ell}>0$ such that $\|\nabla\hat{J}{{}^{MFPG}}(\theta)\|\leq\hat{\ell}~\forall~\theta\in^{d}$. Further, Consider the random variable $W_{k}$, defined as
@@
Using Lemmas˜1 and 2, we can now formally prove Theorem˜1, which we restate for the reader’s convenience.
@@
where $L_{T}$ is the Lipschitz constant of the high-fidelity policy gradient, established in Lemma˜1. Moreover, we recover the corresponding rate for the high-fidelity only policy iterates $\{\theta_{k}^{h}\}_{k\in\mathbb{Z}^{+}}$ by substituting $\rho=0$. Finally, we have that $\lim_{k\to\infty}\mathbb{E}\left[\|\nabla J(\theta_{k}^{MFPG})\|\right]=0$ almost surely.
@@
We establish the global asymptotic convergence of the multi-fidelity policy gradient using the arguments developed in Zhang et al. (2020). To show that $\lim_{k\to\infty}\|\nabla J(\theta_{k}^{MFPG})\|=0$ almost surely, we first establish that $\liminf_{k\to\infty}\|\nabla J(\theta_{k}^{MFPG})\|=0$. Let $J^{*}$ be the global optimum, we have that $W_{k}\leq J^{*}$ by definition of $W_{k}$. From the stochastic ascent property and corresponding bound established in Lemma˜2, we have
@@
which leads us to a contradiction in light of Equation˜6, and thus $\limsup_{k\to\infty}\|\nabla J(\theta_{k}^{MFPG})\|=0$ almost surely. Thus, we get that $\lim_{k\to\infty}\|\nabla J(\theta_{k}^{MFPG})\|=0$.
Appendix C Complete Experimental Results on Off-Dynamics Tasks
Section˜6.3 presents a representative subset of 15 task settings for clarity. Here, we report the complete results across all 39 task settings, spanning diverse types and levels of multi-fidelity dynamics gaps. The overall trends are consistent with those in Section˜6.3. Figures˜12 and 13 summarize the final evaluation return and AUC metrics, respectively, across the full set of off-dynamics tasks.
Figure 12: Final evaluation returns of MFPG versus baselines across the full set of off-dynamics task settings described in Section˜6.1.
Figure 13: AUC of MFPG versus baselines across the full set of off-dynamics task settings described in Section˜6.1.
Appendix D Implementation Details and Hyperparameter Setup

### v4

@@ 7 Conclusions
We present a multi-fidelity policy gradient (MFPG) framework, which mixes a small amount of potentially expensive high-fidelity data with a larger volume of cheap lower-fidelity data to construct unbiased, variance-reduced estimators for on-policy policy gradients. We instantiate this framework through a practical, multi-fidelity variant of the standard REINFORCE algorithm. Theoretically, we show that under standard assumptions, the MFPG estimator guarantees asymptotic convergence of multi-fidelity REINFORCE to locally optimal policies (up to first order optimality conditions) in the target high-fidelity environment of interest. Moreover, whenever the statistical correlation between high- and low-fidelity samples is nontrivial, MFPG REINFORCE achieves faster finite-sample convergence than its single-fidelity counterpart. Empirical evaluation of MFPG in high-fidelity data-scarce regimes across a suite of simulated robotics benchmark tasks with varied multi-fidelity transition dynamics highlights both its reliability and robustness. In our baseline comparison experiments, supported by a 20-seed statistical analysis, for scenarios where low-fidelity data are neutral or beneficial and dynamics gaps are mild to moderate, MFPG is, among the evaluated off-dynamics RL and low-fidelity-only approaches, the only method that consistently achieves statistically significant improvements in mean performance over a baseline trained solely on high-fidelity data. When low-fidelity data become harmful, MFPG exhibits the strongest robustness against performance degradation among the evaluated methods. By contrast, strong off-dynamics RL methods tend to exploit low-fidelity data more aggressively, and can achieve high performance at times, but can also fail substantially more severely. Furthermore, an additional experiment in which the high- and low-fidelity environments are assigned anti-correlated rewards demonstrates that MFPG can remain effective even in the presence of reward misspecification in the low-fidelity environment.
In summary, the proposed framework offers a reliable and robust paradigm for leveraging multi-fidelity data in reinforcement learning, providing promising directions for efficient sim-to-real transfer and principled approaches to managing the trade-off between policy performance and data collection costs.
Limitations & future work.
Future research spans several avenues. First, we study the classic REINFORCE algorithm as a starting point to ease theoretical analysis. Extending the MFPG framework to modern RL algorithms—including actor–critic, off-policy, and offline methods—raises new questions about bias–variance trade-offs that merit deeper investigation. We defer a more detailed discussion of this point to Appendix˜G.
Second, MFPG relies on nontrivial statistical correlation between multi-fidelity samples to achieve variance reduction. Developing systematic techniques to enhance correlation between high- and low-fidelity samples without introducing bias is critical for further performance gains.
Third, our analysis and implementation make several simplifying assumptions, and extending the MFPG framework to more general settings is of great importance. These extensions include (i) integrating data from an arbitrary number of environments, each simulating the target environment with a different fidelity level and cost per sample, (ii) handling multi-fidelity MDPs with different state and action spaces, initial-state distributions, or discount factors, (iii) extending MFPG to infinite-horizon MDPs, and (iv) devising mechanisms to sample correlated trajectories even when low-fidelity simulators do not support resets to arbitrary user-specified initial states.
Finally, applying MFPG to real-world robotic systems represents an important step toward realizing its potential in practical sim-to-real transfer.
@@ Acknowledgments
We thank Haoran Xu, Mustafa Karabag, Brett Barkley, and Jacob Levy for their helpful discussions. This work was supported in part by the National Science Foundation (NSF) under Grants 2409535 and 2214939, in part by the Defense Advanced Research Projects Agency (DARPA) under the Transfer Learning from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) program under grant number HR0011-24-9-0431, and in part by the DEVCOM Army Research Laboratory under Cooperative Agreement Numbers W911NF-23-2-0011 and W911NF-25-2-0021. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Complete MFPG REINFORCE Algorithm
1
Input: Initial policy $\pi_{\theta_{1}}$, value network $V_{\phi}$, multi-fidelity environments $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$, batch size for correlated trajectory samples $N^{h}$, batch size for control variate sample mean estimate $N^{l}$, step sizes $\{\alpha_{k}\}$, maximum number of iterations NN.
2
Output: Trained policy $\pi_{\theta_{N}}$.
3
4for $k=1$ to NN do
5
  // Current policy at iteration $k$: $\pi_{\theta_{k}}(a\mid s)$
6
  Collect $N^{h}$ correlated trajectory samples $\mathcal{D}_{1}=\{\tau_{i}^{h},\tau_{i}^{l}\}_{i=1}^{N^{h}}$ using $\pi_{\theta_{k}}$.
 // See Algorithm 1
7
8  Collect $N^{l}$ additional uncorrelated low-fidelity trajectory samples $\mathcal{D}_{2}=\{\tau_{j}^{l}\}_{j=1}^{N^{l}}$ using $\pi_{\theta_{k}}$.
9  Compute correlated policy gradient random variables using $\mathcal{D}_{1}$:
${\color[rgb]{0,0,0}X^{\pi_{\theta_{k}}}_{\tau^{h},i}}=\frac{1}{T}\sum_{t=0}^{T-1}(G^{h}_{i,t}-V_{\phi}(s^{h}_{i,t}))\,\log{\color[rgb]{0,0,0}\pi_{\theta_{k}}}(a^{h}_{i,t}\mid s^{h}_{i,t}),$
${\color[rgb]{0,0,0}X^{\pi_{\theta_{k}}}_{\tau^{l},i}}=\frac{1}{T}\sum_{t=0}^{T-1}(G^{l}_{i,t}-V_{\phi}(s^{l}_{i,t}))\,\log{\color[rgb]{0,0,0}\pi_{\theta_{k}}}(a^{l}_{i,t}\mid s^{l}_{i,t}).$
10  Estimate the optimal control variate coefficient $\hat{c}_{k}^{*}$ using $\mathcal{D}_{1}$ and Eq.˜4.
11  Estimate the control variate sample mean using $\mathcal{D}_{2}$: ${\color[rgb]{0,0,0}\hat{\mu}^{l}_{k}}=\frac{1}{N^{l}}\sum_{j=1}^{N^{l}}{\color[rgb]{0,0,0}X^{\pi_{\theta_{k}}}_{\tau^{l},j}}$.
12  Construct MFPG estimator: ${\color[rgb]{0,0,0}\nabla_{\theta_{k}}}\mathbb{E}[Z^{{\color[rgb]{0,0,0}\pi_{\theta_{k}}}}]\approx{\color[rgb]{0,0,0}\widehat{g}^{\text{MFPG}}_{k}=}\frac{1}{N^{h}}\sum_{i=1}^{N^{h}}\nabla_{{\color[rgb]{0,0,0}\theta_{k}}}\bigl[{\color[rgb]{0,0,0}X^{\pi_{\theta_{k}}}_{\tau^{h},i}}+{\color[rgb]{0,0,0}\hat{c}_{k}^{*}}({\color[rgb]{0,0,0}X^{\pi_{\theta_{k}}}_{\tau^{l},i}}-{\color[rgb]{0,0,0}\hat{\mu}_{k}^{l}})\bigr]$.
13  Perform policy update $\theta_{k+1}\leftarrow\theta_{k}+\alpha_{k}\,\widehat{g}^{\text{MFPG}}_{k}$.
14  Fit the value function network $V_{\phi}$ using high-fidelity samples in $\mathcal{D}_{1}$.
15return policy $\pi_{{\color[rgb]{0,0,0}\theta_{N}}}$
@@ 9  Compute correlated policy gradient random variables using $\mathcal{D}_{1}$:
Appendix B Proof of Lemma˜1
The proof of Lemma˜1 follows standard control variate arguments (Owen, 2013, Ch.8) and proceeds as follows.
Proof.
We first establish unbiasedness of the control variate estimator for an arbitrary coefficient $c\in\mathbb{R}$. Taking expectation for the defined multi-fidelity control variate random variable over the joint probability space $\Omega$ defined in Section˜4 yields
$\mathbb{E}_{\omega\sim\Omega}[Z^{\pi_{\theta}}(c)]$ $=\mathbb{E}_{\omega\sim\Omega}[X_{\tau^{h}}^{\pi_{\theta}}+c(X_{\tau^{l}}^{\pi_{\theta}}-\mu^{l})]$
$=\mathbb{E}_{\omega\sim\Omega}[X_{\tau^{h}}^{\pi_{\theta}}]+c(\underbrace{\mathbb{E}_{\omega\sim\Omega}[X_{\tau^{l}}^{\pi_{\theta}}]-\mu^{l}}_{=0})$
$=\mathbb{E}_{\omega\sim\Omega}[X_{\tau^{h}}^{\pi_{\theta}}]$
$=\mathbb{E}_{\tau^{h}\sim\mathcal{M}^{h}(\pi_{\theta})}[X_{\tau^{h}}^{\pi_{\theta}}],$
where sampling from the joint probability space $\omega\sim\Omega$ reduces to sampling single-fidelity trajectories $\tau^{h}\sim\mathcal{M}^{h}(\pi_{\theta})$ and $\tau^{l}\sim\mathcal{M}^{l}(\pi_{\theta})$ when evaluating the expectation of the corresponding single-fidelity random variables $X_{\tau^{h}}^{\pi_{\theta}}$ and $X_{\tau^{l}}^{\pi_{\theta}}$, respectively.
Next, we examine the variance of the multi-fidelity control variate estimator:
$\textrm{Var}\bigl(Z^{\pi_{\theta}}(c)\bigr)$ $=\textrm{Var}\bigl(X_{\tau^{h}}^{\pi_{\theta}}+c(X_{\tau^{l}}^{\pi_{\theta}}-\mu^{l})\bigr)$
$=\textrm{Var}\bigl(X_{\tau^{h}}^{\pi_{\theta}}+cX_{\tau^{l}}^{\pi_{\theta}}\bigr)$
$=\textrm{Var}(X_{\tau^{h}}^{\pi_{\theta}})+c^{2}\textrm{Var}(X_{\tau^{l}}^{\pi_{\theta}})+2c\textrm{Cov}(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}}),$
where the second line uses the fact that adding/subtracting a constant does not change the variance of a random variable, and the third line uses the standard formula for the variance of a linear combination of random variables. The variance $\textrm{Var}\bigl(Z^{\pi_{\theta}}(c)\bigr)$ defines a quadratic function in $c$, whose minimizer can be found by setting
$\nabla_{c}\textrm{Var}\bigl(Z^{\pi_{\theta}}(c)\bigr)=0\implies c^{*}=-\frac{\textrm{Cov}(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})}{\textrm{Var}(X_{\tau^{l}}^{\pi_{\theta}})}=-\rho(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})\frac{\sqrt{\textrm{Var}(X_{\tau^{h}}^{\pi_{\theta}})}}{\sqrt{\textrm{Var}(X_{\tau^{l}}^{\pi_{\theta}})}}.$
At the minimizer $c^{*}$, the variance of the multi-fidelity control variate estimator is
$\textrm{Var}\bigl(Z^{\pi_{\theta}}(c^{*})\bigr)$ $=\textrm{Var}(X_{\tau^{h}}^{\pi_{\theta}})+\frac{{\textrm{Cov}^{2}(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})}}{\textrm{Var}(X_{\tau^{l}}^{\pi_{\theta}})}-\frac{2{\textrm{Cov}^{2}(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})}}{\textrm{Var}(X_{\tau^{l}}^{\pi_{\theta}})}$
$=\textrm{Var}(X_{\tau^{h}}^{\pi_{\theta}})-\frac{\rho^{2}(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})\textrm{Var}(X_{\tau^{h}}^{\pi_{\theta}})\textrm{Var}(X_{\tau^{l}}^{\pi_{\theta}})}{\textrm{Var}(X_{\tau^{l}}^{\pi_{\theta}})}$
$=\bigl(1-\rho^{2}(X_{\tau^{h}}^{\pi_{\theta}},X_{\tau^{l}}^{\pi_{\theta}})\bigr)\textrm{Var}(X_{\tau^{h}}^{\pi_{\theta}}).$
Appendix C Proof of Theorem˜1
@@
Thus, for $\theta_{1},\theta_{2}\in\real^{d}$, we have for the high-fidelity only policy gradient,
$\left\|\nabla J(\theta_{1})-\nabla J(\theta_{2})\right\|=\Bigg\|\sum_{t=0}^{T}\sum_{\kappa=0}^{T-t}\gamma^{t+\kappa}\int\Big(\underbrace{r^{h}(s^{h}_{t+\kappa},a^{h}_{t+\kappa})\left[\nabla\log\pi_{\theta_{1}}(a_{t}^{h}\mid s_{t}^{h})-\nabla\log\pi_{\theta_{2}}(a_{t}^{h}\mid s_{t}^{h})\right]p^{h}_{\theta_{1},0:t+\kappa}}_{A}$
$+\underbrace{r^{h}(s^{h}_{t+\kappa},a^{h}_{t+\kappa})\nabla\log\pi_{\theta_{2}}(a_{t}^{h}\mid s_{t}^{h})\left(p^{h}_{\theta_{1},0:t+\kappa}-p^{h}_{\theta_{2},0:t+\kappa}\right)}_{B}\Big)ds^{h}_{0:t+\kappa}da^{h}_{0:t+\kappa}\Bigg\|$
$\leq\sum_{t=0}^{\infty}\sum_{\kappa=0}^{\infty}\gamma^{t+\kappa}\Bigg[\underbrace{\int\Big(\left|r^{h}(s^{h}_{t+\kappa},a^{h}_{t+\kappa})\right|\left\|\nabla\log\pi_{\theta_{1}}(a_{t}^{h}\mid s_{t}^{h})-\nabla\log\pi_{\theta_{2}}(a_{t}^{h}\mid s_{t}^{h})\right\|p^{h}_{\theta_{1},0:t+\kappa}ds^{h}_{0:t+\kappa}da^{h}_{0:t+\kappa}}_{I_{1}}$
$+\underbrace{\int\left|r^{h}(s^{h}_{t+\kappa},a^{h}_{t+\kappa})\right|\left\|\nabla\log\pi_{\theta_{2}}(a_{t}^{h}\mid s_{t}^{h})\right\|\left|p^{h}_{\theta_{1},0:t+\kappa}-p^{h}_{\theta_{2},0:t+\kappa}\right|ds^{h}_{0:t+\kappa}da^{h}_{0:t+\kappa}}_{I_{2}}\Bigg]$ (5)
@@
$\sum_{t=0}^{T}\sum_{\kappa=0}^{T-t}(t+\kappa+1)\gamma^{t+\kappa}$ $=\frac{\left(1-(T+2)^{2}\gamma^{T+1}+(T+3)(T+1)\gamma^{T+2}\right)(1-\gamma)}{(1-\gamma)^{3}}$
$\quad+\frac{2\gamma\left(1-(T+2)\gamma^{T+1}+(T+1)\gamma^{T+2}\right)}{(1-\gamma)^{3}},$
@@
Under Assumptions˜1, 2, 3 and 4, the norm of the Multi-Fidelity Policy Gradient estimator is bounded, i.e., $\exists~\hat{\ell}>0$ such that $\|\nabla\hat{J}{{}^{MFPG}}(\theta)\|\leq\hat{\ell}~\forall~\theta\in\real^{d}$. Further, Consider the random variable $W_{k}$, defined as
@@
Using Lemmas˜2 and 3, we can now formally prove Theorem˜1, which we restate for the reader’s convenience.
@@
where $L_{T}$ is the Lipschitz constant of the high-fidelity policy gradient, established in Lemma˜2. Moreover, we recover the corresponding rate for the high-fidelity only policy iterates $\{\theta_{k}^{h}\}_{k\in\mathbb{Z}^{+}}$ by substituting $\rho=0$. Finally, we have that $\lim_{k\to\infty}\mathbb{E}\left[\|\nabla J(\theta_{k}^{MFPG})\|\right]=0$ almost surely.
@@
We establish the global asymptotic convergence of the multi-fidelity policy gradient using the arguments developed in Zhang et al. (2020). To show that $\lim_{k\to\infty}\|\nabla J(\theta_{k}^{MFPG})\|=0$ almost surely, we first establish that $\liminf_{k\to\infty}\|\nabla J(\theta_{k}^{MFPG})\|=0$. Let $J^{*}$ be the global optimum, we have that $W_{k}\leq J^{*}$ by definition of $W_{k}$. From the stochastic ascent property and corresponding bound established in Lemma˜3, we have
@@
which leads us to a contradiction in light of Equation˜8, and thus $\limsup_{k\to\infty}\|\nabla J(\theta_{k}^{MFPG})\|=0$ almost surely. Thus, we get that $\lim_{k\to\infty}\|\nabla J(\theta_{k}^{MFPG})\|=0$.
Appendix D Additional Ablation Study Results
This section presents additional ablation results that complement Section˜6.5.
Reparameterization trick.
[…]
This pattern is intuitive: when low-fidelity dynamics are close to those of the high-fidelity environment, stochasticity in action selection dominates; reparameterization controls this source of randomness and produces highly correlated rollouts. In the extreme case where the low-fidelity dynamics are perfect (and deterministic), reparameterization makes the low- and high-fidelity rollouts identical, so that $X^{\pi_{\theta}}_{\tau^{h}}$ and $X^{\pi{\theta}}_{\tau^{l}}$ cancel out. The policy gradient estimate then reduces to $\mu^{l}$, which is estimated with abundant (perfect) low-fidelity samples. By contrast, when the dynamics gap is large, discrepancies in transition dynamics dominate, and even with reparameterization, divergence between high- and low-fidelity rollouts persists, resulting in weaker correlation.
Figure 10: Evaluation return curves for MFPG with and without the reparameterization trick versus High-Fidelity Only on Hopper-v3 with 0.5× (left) and 1.2× (right) friction. Solid lines indicate bootstrap point estimates (means), and shaded regions denote two-sided 95% bootstrap confidence intervals. The reparameterization trick proves especially critical in mild-gap settings (e.g., 1.2× friction).
Figure 11: Evaluation return curves for MFPG with and without periodic reconstraining of low-fidelity trajectories versus High-Fidelity Only on Hopper-v3 with 0.5× (left) and 1.2× (right) friction. Solid lines indicate bootstrap point estimates (means), and shaded regions denote two-sided 95% bootstrap confidence intervals. Reconstraining can accelerate learning in some settings—for example, around 0.05M steps in the right panel—but can also harm performance in others, e.g., during the second half of training in the left panel. Accordingly, we do not employ this trick in our evaluations of MFPG.
Variance-bias trade-off.
[…]
Appendix E Additional Experimental Results on Off-Dynamics Tasks
For the 15 main task settings in Section˜6.3 evaluated with 20 random seeds per method, Fig.˜12 reports the mean difference in AUC of each method relative to the High-Fidelity Only baseline (so 0 corresponds to High-Fidelity Only), computed analogously to Fig.˜3. Bars show the bootstrap point estimate of the mean difference, and error bars denote two-sided 95% bootstrap confidence intervals. As a complement to Fig.˜5, Fig.˜13 reports the mean difference in evaluation return between MFPG and the High-Fidelity Only baseline.
Figures˜14 and 15 further show box plots of the final evaluation return and AUC for each method and setting across the 20 runs: the center line indicates the median, the box spans the interquartile range (25th–75th percentiles), and whiskers extend to the 5th and 95th percentiles; outliers beyond the whiskers are not shown.
Figures˜16 and 17 further reports results on 24 additional task settings, each evaluated with 5 random seeds per method. Due to the small number of seeds, these additional results are primarily used to illustrate the consistency of qualitative trends (e.g., the prevalence of performance collapses), and we do not draw formal statistical conclusions from them. Instead of reporting full distributions, in these figures, bar heights indicate medians and error bars denote the minimum and maximum values across seeds.
In Figs.˜16 and 17, kinematic variations are categorized as easy, medium, or hard; each setting constrains a specific joint to a reduced range of motion to mimic actuator damage. For kinematic shifts, we consider cases where the agent’s thigh joint is impaired (leg joint for Hopper-v3, since the original benchmark does not include a thigh joint shift for Hopper).
[…]
Figure 12: Mean difference in AUC of each method relative to the High-Fidelity Only baseline (0 corresponds to High-Fidelity Only) across the 15 main task settings in Section˜6.3, computed analogously to Fig.˜3. Bars show the bootstrap point estimate (mean), and error bars denote two-sided 95% bootstrap confidence intervals.
Figure 13: Mean difference in evaluation return of MFPG relative to the High-Fidelity Only baseline (0 indicates no difference) for the settings in Fig.˜5. Solid lines show the bootstrap mean estimate, and shaded regions denote two-sided 95% bootstrap confidence intervals.
Figure 14: Box plots of the final evaluation return for each method across the 15 main task settings in Section˜6.3. The center line indicates the median, the box spans the interquartile range (25th–75th percentiles), and whiskers extend to the 5th and 95th percentiles; outliers beyond the whiskers are not shown.
Figure 15: Box plots of the AUC for each method across the 15 main task settings in Section˜6.3. The center line indicates the median, the box spans the interquartile range (25th–75th percentiles), and whiskers extend to the 5th and 95th percentiles; outliers beyond the whiskers are not shown.
Figure 16: Final evaluation return of each method across an additional 24 task settings. Bar heights indicate medians, and error bars denote the minimum and maximum values across seeds.
Figure 17: AUC of each method across an additional 24 task settings. Bar heights indicate medians, and error bars denote the minimum and maximum values across seeds.
Appendix F Implementation Details and Hyperparameter Setup